A semi-annual coupon bond has MacD of 27.0 years, yield-to-maturity of 4.0%, and price of $1200.00. What is its DV01? Answer in dollars, rounded to three decimal places.
Answer #1

Macaulay Duration = 27 years

Modified Duration = 27/1.04 = 25.96 years

DV01 = Modified Duration(Price)/10,000

DV01 = 25.96(1,200)/10,000

DV01 = $3.115